Calculate the volume, in milliliters, of a 0.490 M KOH solution that should be added to 4.250 g of HEPES (MW = 238.306 g/mol, pKa = 7.56) to give a pH of 7.94.

KOH
Answer #1

Ans. 39.3ml

Moles of HPES used & graus of UPES. Molar Mass of UPES 9.2509 0.0178 Moles 238.306 g/mol LA) 7.94 = 7.56 + oso sag just As we

So, at equilibrium [A-7 = 2.40 [4x ] (from 0 2 ) X = 27.40 10.0178-x) 3.40 X = 2.40 80.0118 X = 0.0126 Moles of A-are = 0.012